OpenStudy (anonymous):

Use the quadratic formula to solve the equation. x2 – 7x – 6 = 0 @ChmE

OpenStudy (anonymous):

\[3x2-2x-1=0\]

OpenStudy (anonymous):

for the first use the quadratic formula, find a, b, and c. Then set up a system of equations, solve with substitution or elimination. I got x=-1 and x=-6 for x2-7x-6=0.

OpenStudy (anonymous):

Rajabu???

OpenStudy (anonymous):

@Rajabu i got to go soon if u need more help say so now!!!

OpenStudy (anonymous):

x^2-7x +6 x^2 -6x -x + 6 x(x-6) -(x-6) (x-1)(x-6) (x-1)(x-6) = 0 x-6 = 0 or x-1 = 0 or x=6 or x = 1